If the angles of a triangle are in the ratio of 2:3:5,what kind of triangle is it?
Answer by checkley77(12844)

2N+3N+5N=10N
10N=180
N=180/10
N=18
2*18=36
3*18=54
5*18=90
ANSWER = A RIGHT TRIANGLE WITH 2 ANGLES OF 36 & 54 DEGREES.